OpenStudy (anonymous):

MEDAL AND FAN Using the completing-the-square method, rewrite f(x) = x2 − 6x + 2 in vertex form. f(x) = (x − 3)2 f(x) = (x − 3)2 + 2 f(x) = (x − 3)2 − 7 f(x) = (x − 3)2 + 9

OpenStudy (jango_in_dtown):

x2 − 6x + 2=x^2-6x+9-7=(x-3)^2-7